# Stochastic gradient descent estimation of generalized matrix factorization models with application to single-cell RNA sequencing data

### v1

@@ 3.2.2 Block-wise adaptive stochastic gradient descent
Orthogonalize $\hat{\mathbf{U}}$ and $\hat{\mathbf{V}}$;
@@
Figure 1: Graphical representation of the stochastic gradient updates employed at the $t$th iteration of Algorithm 2. First column: generalized matrix factorization model. Second column: matrix operations employed to obtain the gradients of the penalized log-likelihood. Third column: adaptive gradient step for updating the parameters. The coloured cells highlight the sub-sampled data used for the $t$th update. Empty cells represent the data not used at the $t$th update. To save space, the calculations of the second-order differentials and the gradient smoothing are not displayed here.
We refer to (14) as local or partial updates since they just modify the rows of $\mathbf{U}$ and $\mathbf{V}$ corresponding to the minibatch block of indices $B=I\times J$. On the opposite, we refer to updates (10) and (13) as global updating rules.
The choice between alternated least squares (Kidziński et al.,, 2022; Wang and Carvalho,, 2023), exact quasi-Newton (Algorithm 1; Kidziński et al.,, 2022) and the proposed block-wise stochastic gradient descent (Algorithm 2, Figure 1) methods is up to the researcher and, in principle, depends on the dimension of the problem under study and on the available computational resources. In general, algorithms using a higher amount of information are more stable, reliable and accurate; however, they tend to scale poorly in high-dimensional settings. On the other hand, cheaper methods that use less information scale well in big-data problems at the cost of a lower level of precision.
3.3 Missing value imputation
Whenever the input data matrix is only partially complete and we need to maximize (6) over the set of observed data, it is necessary to properly handle the missing values during the estimation process. To this end, we rely on the general framework proposed by Cai et al., (2010) and Mazumder et al., (2010), and later used by Kidziński et al., (2022). Intuitively, it prescribes imputing the missing data entries during the optimization by updating them at each iteration using the most recent prediction of those values.
To formally describe this imputation scheme, we define $P_{\Omega}(\mathbf{Y})$ as the projection of $\mathbf{Y}$ onto its observed entries, say $P_{\Omega}(\mathbf{Y})_{ij}=y_{ij}$ if $(i,j)\in\Omega$ and 00 elsewhere. Similarly, we define $P_{\Omega}^{\perp}(\mathbf{Y})$ as the complementary projection of $P_{\Omega}(\mathbf{Y})$, such that $\mathbf{Y}=P_{\Omega}(\mathbf{Y})+P_{\Omega}^{\perp}(\mathbf{Y})$. Then, in missing value situations, Algorithm 1 and 2 can be adapted by replacing the static (incomplete) data matrix $\mathbf{Y}$ with its completed version $\mathbf{Y}^{t}$ obtained at iteration $t$. And, at the end of each iteration, we introduce the imputation step as
$\mathbf{Y}^{t+1}\leftarrow P_{\Omega}(\mathbf{Y})+P_{\Omega}^{\perp}(\boldsymbol{\mu}^{t+1}).$
As discussed by Mazumder et al., (2010), under the squared-error loss, this online imputation strategy corresponds to an expectation-maximization algorithm, where the missing values are treated as latent variables to be estimated along all the other parameters in a coordinate-wise fashion. In our case, where general exponential family likelihoods are considered and no explicit solutions are available, we can interpret the online estimation of the missing values as an approximated version of the expectation-maximization algorithm.
@@ 3.4 Parameter initialization
As is common for non-convex optimization problems, the performance of Algorithm 2 is strongly influenced by the initial values used to start the optimization cycle. Random initialization may increase the likelihood of convergence to sub-optimal local minima or unstable saddle points. Also, the convergence speed of the algorithm strongly depends on the initial parameter choice. For these reasons, it is fundamental to design good initialization strategies to improve the behaviour of the algorithm and the quality of the final estimates. Here, we propose a deterministic initialization procedure which is compliant with the model assumptions and provide a good approximation of the model at the optimum.
@@
We leverage the conditional GLM structure of model (3) to estimate the regression parameters, $\boldsymbol{\beta}_{j:}$’s and $\boldsymbol{\gamma}_{i:}$’s, solving a sequence of independent likelihood maximization problems. We denote the resulting initial values as $\tilde{\boldsymbol{\beta}}_{j:}$’s and $\tilde{\boldsymbol{\gamma}}_{i:}$’s. Specifically, at the first stage, we estimate the column-specific regression parameters, $\boldsymbol{\beta}_{j:}$’s, fitting $m$ GLMs via standard numerical solvers, e.g., Fisher scoring. Next, we follow the same strategy to estimate the row-specific regression parameters, $\boldsymbol{\gamma}_{i:}$’s, conditionally on the estimated column-specific effects, $\mathbf{x}_{i:}^{\top}\tilde{\boldsymbol{\beta}}_{j:}$, which are used as offset terms. To extract the initial latent score matrix, $\mathbf{U}$, we follow the null-residual strategy proposed by Townes et al., (2019), which consists of computing the first $d$ left eigenvectors of a properly defined residual matrix of the model. Common specifications for such a matrix follow the standard definition of either deviance or Pearson residuals, respectively $r_{ij}^{\textsc{d}}=\textrm{{sign}}(y_{ij}-\mu_{ij})\sqrt{D(y_{ij},\mu_{ij})}$ and $r_{ij}^{\textsc{p}}=(y_{ij}-\mu_{ij})/\sqrt{\nu(\mu_{ij})}$, Here, the conditional mean of $y_{ij}$ can be approximated using the regression effects estimated at the first two stages of the initialization, say $\tilde{\mu}_{ij}=g^{-1}(\mathbf{x}_{i:}^{\top}\tilde{\boldsymbol{\beta}}_{j:}+\tilde{\boldsymbol{\gamma}}_{i:}^{\top}\mathbf{z}_{j:})$. Finally, we compute the initial loading matrix, $\mathbf{V}$, fitting $m$ column-specific GLMs using the matrix of latent scores $\tilde{\mathbf{U}}$ as it was a fixed design matrix and including an offset term in the linear predictor to account for both row- and column-specific regression effects, say $\mathbf{x}_{i:}^{\top}\tilde{\boldsymbol{\beta}}_{j:}+\tilde{\boldsymbol{\gamma}}_{i:}^{\top}\mathbf{z}_{j:}$.
Algorithm 3 provides a pseudo-code description of the proposed initialization procedure, henceforth referred to as GLM–SVD. Using standard numerical solvers to obtain the GLM solutions and the incomplete singular value decomposition, the overall computational complexity of the initialization is of order
[…]
It is worth noticing that each step of the procedure involving GLM fitting is highly parallelizable, where each independent subroutine has at most a computational complexity of $O(k^{3}+k^{2}n)$ where $k=\max(p,q,d)$. Improved rates can be obtained by employing alternative solvers, such as coordinate ascent (Friedman et al.,, 2007, 2010) or stochastic gradient methods (Toulis and Airoldi,, 2015; Tran et al.,, 2015).
Memory allocation
$\tilde{\boldsymbol{\eta}}\leftarrow\mathbf{0}_{n\times m}$;  $\tilde{\mathbf{B}}\leftarrow\mathbf{0}_{m\times p}$;  $\tilde{\boldsymbol{\Gamma}}\leftarrow\mathbf{0}_{n\times q}$;
Initialization of $\mathbf{B}$
for $j=1,\dots,m$ do
       $\tilde{\boldsymbol{\beta}}_{j:}\leftarrow\texttt{glm.fit}(\texttt{x}=\mathbf{X},\;\texttt{y}=\mathbf{y}_{:j},\;\texttt{offset}=\tilde{\boldsymbol{\eta}}_{:j},\;\texttt{family}=\texttt{EF})$; $O(p^{3}+p^{2}n)$
      $\tilde{\boldsymbol{\eta}}_{:j}\leftarrow\tilde{\boldsymbol{\eta}}_{:j}+\mathbf{X}\tilde{\boldsymbol{\beta}}_{:j}$; $O(pn+n)$
Initialization of $\boldsymbol{\Gamma}$
for $i=1,\dots,n$ do
       $\tilde{\boldsymbol{\gamma}}_{i:}\leftarrow\texttt{glm.fit}(\texttt{x}=\mathbf{Z},\;\texttt{y}=\mathbf{y}_{i:},\;\texttt{offset}=\tilde{\boldsymbol{\eta}}_{i:},\;\texttt{family}=\texttt{EF})$; $O(q^{3}+q^{2}m)$
      $\tilde{\boldsymbol{\eta}}_{i:}\leftarrow\tilde{\boldsymbol{\eta}}_{i:}+\mathbf{Z}\tilde{\boldsymbol{\gamma}}_{i:}$; $O(qm+m)$
Initialization of $\mathbf{U}$
$\tilde{\boldsymbol{\mu}}\leftarrow g^{-1}(\tilde{\boldsymbol{\eta}})$;  $\tilde{\mathbf{R}}\leftarrow\texttt{resid}(\mathbf{Y},\tilde{\boldsymbol{\mu}})$;  $\tilde{\mathbf{U}},\tilde{\mathbf{D}},\tilde{\mathbf{V}}\leftarrow\texttt{svd}(\tilde{\mathbf{R}})$; $O(dnm+nm)$
Initialization of $\mathbf{V}$
for $j=1,\dots,m$ do
       $\tilde{\mathbf{v}}_{:j}\leftarrow\texttt{glm.fit}(\texttt{x}=\tilde{\mathbf{U}},\;\texttt{y}=\mathbf{y}_{:j},\;\texttt{offset}=\tilde{\boldsymbol{\eta}}_{:j},\;\texttt{family}=\texttt{EF})$; $O(d^{3}+d^{2}n)$
      $\tilde{\boldsymbol{\eta}}_{:j}\leftarrow\tilde{\boldsymbol{\eta}}_{:j}+\tilde{\mathbf{U}}\tilde{\mathbf{v}}_{:j}$; $O(dn+n)$
Algorithm 3 Pseudo-code description of the GLM–SVD initialization algorithm described in Section 3.4. On the right, we report the computation complexity of each step. The compact notation glm.fit() and svd() follow the naming conventions used in the corresponding R functions, assuming that a suitable model family, say EF, has been selected.
As an alternative strategy for high-dimensional settings, we propose to approximate the GLM estimates using a misspecified Gaussian model over a link-based transformation of the response matrix, say $y_{\varepsilon,ij}=g_{\varepsilon}(y_{ij})$. We adopt the perturbed link function $g_{\varepsilon}(\cdot)$ in place of $g(\cdot)$ to prevent the occurrence of infinite values when evaluating the link transformation, a common adjustment for binary and count data. Then, we can initialize the regression coefficient and the loading matrices solving sequentially the following least squares problems: $\min_{\boldsymbol{\beta}}\|\mathbf{Y}_{\varepsilon}-\mathbf{X}\mathbf{B}^{\top}\|^{2}$, $\min_{\boldsymbol{\gamma}}\|\mathbf{Y}_{\varepsilon}-\mathbf{X}\tilde{\mathbf{B}}^{\top}-\boldsymbol{\Gamma}\mathbf{Z}^{\top}\|^{2}$ and $\min_{\mathbf{v}}\|\mathbf{Y}_{\varepsilon}-\mathbf{X}\tilde{\mathbf{B}}^{\top}-\tilde{\boldsymbol{\Gamma}}\mathbf{Z}^{\top}-\tilde{\mathbf{U}}\mathbf{V}^{\top}\|^{2}$. We can compute the factor scores $\tilde{\mathbf{U}}$ using the same residual decomposition discussed so far. Algorithm 4 provides a pseudo-code description of this second initialization procedure, which reduces the computation complexity of the initialization to
[…]
By doing so, the cubic and quadratic terms in $p$, $q$ and $d$ are no longer multiplied by an external factor dependent on $n$ or $m$. As a consequence, the overall computation complexity is $O\{(p^{3}+q^{3}+d^{3})+(p^{2}n+q^{2}m+d^{2}n)+(p+q+d)nm\}$. Hence, in high-dimensional settings with $\max(p,q,d)\ll\min(n,m)$, the computational cost is dominated by the leading term $(p+q+d)nm$. We refer to such an approximation as OLS–SVD.
Data transformation
$\mathbf{Y}_{\varepsilon}\leftarrow g_{\varepsilon}(\mathbf{Y})$;  $\tilde{\boldsymbol{\eta}}\leftarrow\mathbf{0}_{n\times m}$; $O(nm)$
Initialization of $\mathbf{B}$
$\tilde{\mathbf{B}}\leftarrow(\mathbf{Y}_{\varepsilon}-\tilde{\boldsymbol{\eta}})^{\top}\mathbf{X}\,(\mathbf{X}^{\top}\mathbf{X})^{-1}$;  $\tilde{\boldsymbol{\eta}}\leftarrow\tilde{\boldsymbol{\eta}}+\mathbf{X}\tilde{\mathbf{B}}^{\top}$; $O(p^{3}+p^{2}n+pnm)$
Initialization of $\boldsymbol{\Gamma}$
$\tilde{\boldsymbol{\Gamma}}\leftarrow(\mathbf{Y}_{\varepsilon}-\tilde{\boldsymbol{\eta}})\,\mathbf{Z}\,(\mathbf{Z}^{\top}\mathbf{Z})^{-1}$;  $\tilde{\boldsymbol{\eta}}\leftarrow\tilde{\boldsymbol{\eta}}+\tilde{\mathbf{A}}\mathbf{Z}^{\top}$; $O(q^{3}+q^{2}n+qnm)$
Initialization of $\mathbf{U}$
$\tilde{\boldsymbol{\mu}}\leftarrow g^{-1}(\tilde{\boldsymbol{\eta}})$;  $\tilde{\mathbf{R}}\leftarrow\texttt{resid}(\mathbf{Y},\tilde{\boldsymbol{\mu}})$;  $\tilde{\mathbf{U}},\tilde{\mathbf{D}},\tilde{\mathbf{V}}\leftarrow\texttt{svd}(\tilde{\mathbf{R}})$; $O(dnm+nm)$
Initialization of $\mathbf{V}$
$\tilde{\mathbf{V}}\leftarrow(\mathbf{Y}_{\varepsilon}-\tilde{\boldsymbol{\eta}})^{\top}\tilde{\mathbf{U}}\,(\tilde{\mathbf{U}}^{\top}\tilde{\mathbf{U}})^{-1}$; $O(d^{3}+d^{2}n+dnm)$
Algorithm 4 Pseudo-code description of the OLS–SVD initialization algorithm described in Section 3.4. On the right, we report the computation complexity of each step.
While inspired by the null-residual strategy of Townes et al., (2019), our proposed initialization methods diverge from their approach in four key aspects. Firstly, our method is exclusively employed as an initialization strategy and is not utilized as an approximation to the final estimates. Secondly, our proposal is adaptable to any exponential family model, unlike the original approximation, which was tailored solely for count observations and lacked direct applicability to binary or continuous data. Thirdly, our proposal explicitly addresses covariate-dependent sampling schemes; in contrast, the original implementation by Townes and co-authors relied on an independent approximation of the null-residuals based on the properties of the contingency table representation of the response matrix. Lastly, our regression-based strategy for computing the loading matrix ensures that the solution approximately satisfies a proper conditional estimating equation, whereas the left eigenvectors employed in the null-residual approach do not guarantee the nullification of a proper score equation.
4 Model selection
In the GMF formulation detailed in Section 2, the model complexity is mainly controlled by the rank of the matrix factorization, $d$. Rank selection in matrix factorization problems is a complex research problem far from being solved. As for many other methods in statistics, the optimal selection of the factorization rank needs a careful balance between model complexity and goodness of fit to avoid both under- and over-fitting issues. In the matrix factorization literature, four main approaches can be adopted for the rank selection.
@@ 4 Model selection
Eigenvalue thresholding criteria. The first method is to rely on the so-called eigenvalue thresholding, or elbow, criterion. It involves analyzing the singular values, in descending order, of a sufficiently high-rank matrix decomposition. The point at which we can detect a significant decrease in the rate of change of the singular values suggests the optimal number of factors to retain. This point signifies the balance between capturing sufficient variance in the data and avoiding overfitting. Thanks to its simplicity, computational efficiency, and the significant amount of empirical and theoretical support, eigenvalue thresholding methods gained wide popularity for conventional rank selection problems. See, e.g., the work of Onatski, (2010), Fan et al., (2022) and Wang and Carvalho, (2023). Such a class of selection criteria favors a compact low-rank representation of the signal and is often used when it is of interest to identify the principal modes of variation in the data for interpretation purposes.
Information-based criteria. The second approach is to select the unknown rank by minimizing some appropriate information criteria, such as the Akaike (AIC) or the Bayesian (BIC) information. The information, in these cases, is defined as an adaptation measure which tries to balance an in-sample goodness-of-fit summary, the negative log-likelihood function, and a complexity score which typically depends on the sample size and on the number of unknown parameters to be estimated. For instance, in the proposed GMF model, the sample size corresponds to the number of available entries in the data matrix, $nm$ for complete data, and the dimension of the parameter space is $nq+mp+(n+m)d+1$. Such a model selection criterion can be quite expensive since it requires the estimation of several models, one for each rank to be tested. However, computations can be accelerated by considering clever warm start strategies within a forward/backward stagewise estimation approach. Unfortunately, there are no guarantees that different information criteria give similar solutions, at least for finite samples.
Out-of-sample error criteria. The third rank determination method proposed in the literature leverages the concept of out-of-sample error minimization. An estimate of the out-of-sample error for matrix factorization problems can be obtained by holding out a bunch of entries of the data matrix during the estimation and then looking at the reconstruction error either after or even during the optimization process. The same strategy can be performed repeatedly within a cross-validation procedure to obtain a more reliable estimate of the generalization error. Then, having an error measure for a grid of ranks, we can pick the value minimizing the error. Such an approach is very popular in the machine-learning community for evaluating the performances of recommendation systems, collaborative filtering and matrix completion models, where the main goal is to obtain an accurate reconstruction of the unobserved entries of the response matrix; see, e.g., Koren et al., (2009) and Owen and Perry, (2009) for an extensive statistically flavoured discussion. Rank selection based on error minimization is intuitive, general and fairly robust with respect to the optimization method used for the estimation. However, it might become extremely expensive in terms of computational resources and is highly sensitive to the selection method used for choosing the hold-out set.
Implicit selection via regularizarion. An additional mechanism to control model complexity in the GMF framework is to tune the regularization parameter $\lambda$ in (6), which controls the amount of shrinkage enforced over the latent matrices $\mathbf{U}$ and $\mathbf{V}$. Frobenius regularization is often introduced for numerical stability issues, but it is also intrinsically connected to the rank determination problem. Indeed, for any $\lambda>0$, the Frobenious norm penalty implicitly shrinks the singular values of the matrix factorization toward zero encouraging compact low-rank representations of the signal. This spectral penalization effect and its connection with the nuclear norm has been extensively discussed in, e.g., Witten et al., (2009), Mazumder et al., (2010) and Kidziński et al., (2022). In particular, Kidziński et al., (2022) proposed to implicitly select the optimal number of latent factors in GMF problems by estimating an over-parametrized factorization with fixed rank $d\propto\sqrt{m}$ over a fine grid of $\lambda$-values and minimizing some appropriate error criterion over $\lambda$. The estimation can be efficiently performed using a pathwise optimization scheme with warm start initialization (Friedman et al.,, 2007). Still, it requires the estimation of several over-parametrized GMF models on a fine grid of regularization parameters, which could be quite expensive.
Generally, there is a lack of consensus in the literature regarding the optimal rank selection method, with the choice often contingent upon the objective of the analysis. In exploratory studies aimed at visually uncovering latent structures and interpreting them, eigenvalue-based methods are commonly employed, favouring more conservative approximations. Conversely, for prediction or reconstruction tasks, methods relying on information criteria and out-of-sample error are typically preferred.
In our study, we focus on single-cell RNA sequencing, in which matrix factorization plays a crucial role as a preprocessing tool to remove noise and reduce the dimensionality of the problem. Subsequently, downstream analysis steps, typically in the form of clustering or pseudotime inference, are employed to glean insightful biological information from the latent score and loading matrices. In this context, we advocate for eigenvalue-based rank selection methods, akin to those explored by Wang and Carvalho, (2023). Nevertheless, the computational advantages offered by our stochastic optimization approach could also substantially benefit prediction scenarios, particularly when multiple high-dimensional models need to be fitted to minimize appropriate error criteria across a fine grid of tuning parameters.

### v2

@@ 4. Compute the smoothed gradients and update $\mathbf{U}$
5. Orthogonalize $\hat{\mathbf{U}}$ and $\hat{\mathbf{V}}$;
@@
If the dispersion parameter $\phi$ is unknown and has to be estimated from the data, we can also adopt a smoothed stochastic estimator obtained through exponential averaging. More details are provided in Appendix B in the Supplementary Material.
Whenever the input data matrix is only partially complete, it is necessary to properly handle the missing values during the estimation process. To this end, we rely on the general framework proposed by Cai et al., (2010) and Mazumder et al., (2010), and later used by Kidziński et al., (2022). This prescribes imputing the missing data entries during the optimization by updating them at each iteration using the most recent prediction of those values. Algorithm 1 can thus be adapted by replacing the static (incomplete) minibatch matrix $\mathbf{Y}_{\scriptscriptstyle{B}}$ with its completed version $\mathbf{Y}_{\scriptscriptstyle{B}}^{t}$ obtained at iteration $t$ after imputation. At the beginning of each iteration, we then introduce the imputation step as $y_{ij}^{t}\leftarrow\mu_{ij}^{t}$ if $(i,j)\in B\cap\Omega^{c}$, where $\Omega^{c}$ is the complement of $\Omega$.
We refer to (14) as local or partial updates since they just modify the rows of $\mathbf{U}$ and $\mathbf{V}$ corresponding to the minibatch block of indices $B=I\times J$. On the contrary, we refer to updates (10) and (13) as global updating rules. The choice between alternated least squares (Kidziński et al.,, 2022; Wang and Carvalho,, 2023), exact quasi-Newton (Kidziński et al.,, 2022) and the proposed adaptive stochastic gradient descent (Algorithm 1, Figure 1) methods is up to the researcher and, in principle, depends on the dimension of the problem under study and on the available computational resources. In general, algorithms using a higher amount of information are more stable and accurate; however, they tend to scale poorly in high-dimensional settings and often get stuck on suboptimal optima. On the other hand, cheaper stochastic methods using less information scale well in big-data problems at the cost of a lower level of precision.
4 Additional computational aspects
@@ 4.1 Parameter initialization
As is common in non-convex optimization, the performance of Algorithm 1 heavily depends on its initialization. Random starting values can slow convergence and increase the risk of getting stuck in poor local minima or unstable saddle points. To improve both performance and accuracy, we adopt a structured initialization strategy that leverages the conditional GLM formulation of model (3), and is inspired by the initialization approaches employed in Risso et al., (2018), Townes et al., (2019), and Kidziński et al., (2022).
@@
We first estimate the column-specific regression parameters, $\boldsymbol{\beta}_{j:}$’s, by fitting $m$ separate GLMs. Then, conditionally on the estimated offsets $\mathbf{x}_{i:}^{\top}\tilde{\boldsymbol{\beta}}_{j:}$, we estimate the row-specific parameters, $\boldsymbol{\gamma}_{i:}$’s, using the same strategy, and fitting $n$ separate GLMs. The resulting estimates, $\tilde{\boldsymbol{\beta}}_{j:}$ and $\tilde{\boldsymbol{\gamma}}_{i:}$, initialize the regression effects. To initialize the latent scores, $\mathbf{u}_{i:}$’s, we adopt the null-residual method of Townes et al., (2019), extracting the first $d$ left eigenvectors of a residual matrix based on either deviance or Pearson residuals: $r_{ij}^{\textsc{d}}=\textrm{{sign}}(y_{ij}-\mu_{ij})\sqrt{D(y_{ij},\mu_{ij})}$ and $r_{ij}^{\textsc{p}}=(y_{ij}-\mu_{ij})/\sqrt{\nu(\mu_{ij})}$, where the mean $\mu_{ij}$ is approximated as $\tilde{\mu}_{ij}=g^{-1}(\mathbf{x}_{i:}^{\top}\tilde{\boldsymbol{\beta}}_{j:}+\tilde{\boldsymbol{\gamma}}_{i:}^{\top}\mathbf{z}_{j:})$. Finally, we estimate the loadings, $\mathbf{v}_{j:}$’s, by fitting $m$ column-specific GLMs with the latent scores as fixed design matrix and an offset term $\mathbf{x}_{i:}^{\top}\tilde{\boldsymbol{\beta}}_{j:}+\tilde{\boldsymbol{\gamma}}_{i:}^{\top}\mathbf{z}_{j:}$ in the predictor. Using standard solvers for GLMs and singular value decomposition, the overall computational cost is
$m\,\underbrace{O(p^{3}+p^{2}n)}_{\text{$\textrm{{B}}$--inner cycle}}+\;n\,\underbrace{O(q^{3}+q^{2}m)}_{\text{$\Gamma$--inner cycle}}+\;\underbrace{O((p+q+d)nm)}_{\text{$\textrm{{U}}$--residual SVD}}+\;m\,\underbrace{O(d^{3}+d^{2}n)}_{\text{$\textrm{{V}}$--inner cycle}}.$
All GLM-fitting steps are highly parallelizable, with per-task complexity at most $O(k^{3}+k^{2}n)$, where $k=\max(p,q,d)$. In high-dimensional settings, one may approximate the GLM fits via ordinary least squares on a link-transformed response, reducing complexity to:
$\underbrace{O(p^{3}+p^{2}n+pnm)}_{\text{$\textrm{{B}}$--least squares}}+\;\underbrace{O(q^{3}+q^{2}m+qnm)}_{\text{$\Gamma$--least squares}}+\;\underbrace{O((p+q+d)nm)}_{\text{$\textrm{{U}},\textrm{{V}}$--residual SVD}}.$
Although inspired by the null-residual strategy of Townes et al., (2019), our initialization differs in four key aspects: (i) it serves solely for initialization, not final estimation; (ii) it generalizes to any exponential family, not just count data; (iii) it accounts for covariate-dependent sampling; (iv) it produces loadings satisfying approximate estimating equations, unlike the original purely spectral null-residual approach.
4.2 Model selection
@@
In the GMF formulation detailed in Section 2, the model complexity is mainly controlled by the rank of the matrix factorization, $d$. The optimal selection of the factorization rank needs a careful balance between model complexity and goodness of fit to avoid both under- and over-fitting issues. In the matrix factorization and factor model literature, a popular class of rank selection methods is the so-called spectral thresholding, or elbow, approach. This consists of analyzing the singular values, in descending order, of a sufficiently high-rank matrix decomposition to detect a significant decrease in the rate of change of the singular values, which suggests the optimal number of factors to retain. Thanks to its simplicity, computational efficiency, and the significant amount of empirical and theoretical support, spectral thresholding methods gained wide popularity for rank selection problems. See, e.g., the work of Onatski, (2010), Fan et al., (2022), Wang and Carvalho, (2023), Liu et al., (2023), Nie et al., (2024), and Liu and Zhong, (2024). This class of selection criteria favors a compact low-rank representation of the signal and is often used when it is of interest to identify the principal modes of variation in the data for interpretation purposes.
Another popular rank determination method proposed in the literature leverages the concept of out-of-sample error minimization (Mazumder et al.,, 2010; Kidziński et al.,, 2022). An estimate of the out-of-sample error for matrix factorization problems can be obtained either by using information-based metrics, such as the Akaike (AIC) or the Bayesian (BIC) information criteria, or prediction error measures. The latter may be computed by removing some entries of the data matrix during the estimation process substituting them with missing values and then evaluating the reconstruction error on the hold-out set of data. The same strategy can be performed repeatedly within a cross-validation procedure to obtain a more reliable estimate of the generalization error. The minimizer of such an error measure over a fairly large grid of prespecified matrix ranks is the selected latent dimension.
Rank selection based on error minimization is intuitive, general, and robust with respect to the estimation method. However, its application to high-dimensional data is hindered by its extremely high computational cost, due to the need for multiple estimations of possibly over-parametrized models. Moreover, many state-of-the-art methods proposed in the literature and implemented in standard software packages do not handle missing values directly, thus preventing robust evaluation of out-of-sample error prior to missing value imputation.
In the high-dimensional setting, spectral thresholding methods, akin to those explored by Wang and Carvalho, (2023), Nie et al., (2024), and Liu and Zhong, (2024), gained increasing popularity, thanks to their ease of implementation, computational efficiency and connection with standard scree-plot analysis of PCA.Yet, these are not often used in practice in the omics literature, where practitioners often rely on software default values, such as 10 or 50.
Thanks to its ability of handling missing values, the intrinsic scalability of the proposed stochastic optimization method, and a convenient warm-start initialization strategy (Friedman et al.,, 2007, 2010), our approach enables for the first time in the omics literature the systematic selection of the number of latent factors using well grounded criteria based on cross-validation out-of-sample error and spectral thresholding.